# Adapting LLaMA Decoder to Vision Transformer

### v2

@@ 3.7 Analysis of causal Self-attention
Attention map rank. We examine the representation learning power of causal attention through a spectrum analysis. Following wang2020linformer ; shu2021adder , we perform singular value decomposition on the attention maps of the pre-trained ViT-T/16 and iLLaMA-T/16 models. Next, we sort the singular values and plot a curve illustrating the relationship between the cumulative normalized singular values and matrix indices. The results are conducted using $30$ images randomly selected from the ImageNet-1K validation set. As shown in Figure 4, the curve of ViT exhibits concave function characteristics, while the curve of iLLaMA is close to a linear function, indicating a more uniform distribution of singular values in iLLaMA’s attention map. Approximating the matrix rank by the index at which the cumulative normalized singular value reaches 0.8, we observe that the index value of iLLaMA is about 48 higher than that of ViT ($\sim$129-th v.s. $\sim$81-th). Under such premise, compared to ViT, the attention map of iLLaMA can be approximated with a certain error by a higher-rank matrix. Accordingly, the rank of the attention map may affect the expressive capabilities of the learned representations dong2021attention , suggesting that the causal self-attention in iLLaMA has the potential to learn complex visual representations, as empirically demonstrated in Section 4.2. Detailed results for different layers and heads are provided in Appendix D.
@@ 4.1 Ablation Study
Influence of soft mask for training and testing loss. A common challenge deep neural networks encounter is underfitting liu2023dropout , often characterized by the difficulty in continuously reducing training loss during model training and resulting in unsatisfactory test accuracy. We compare the training and testing losses of the iLLaMA-T/16 and iLLaMA-B/16 models with and without the use of the soft mask strategy, as shown in Table 4. We observe that soft mask can reduce training loss in both regimes, mitigating potential underfitting issue and reducing testing loss.
@@ 4.2 Comparison with Recent Architectures on ImageNet-1K Classification
We conducted experiments on the ImageNet-1K deng2009imagenet benchmark with different model sizes (i.e., iLLaMA-T/S/B/L) to examine its scalability. Detailed architecture configurations are in Appendix A.
ImageNet-1K training. ImageNet-1K dataset contains 1281167 training images and 50000 testing images from 1000 classes. We train iLLaMA-T/S/B on ImageNet-1K for 300 epochs with AdamW optimizer loshchilov2017decoupled and a batch size of 4096. The ImageNet-1K trained iLLaMA-T/B models are, in fact, the outcome of the explorations completed in Section 3.6. For the settings of soft mask schedule, cutoff epochs, and learning rate warmup epochs, we tune slightly for the iLLaMA-S model.
ImageNet-21K pre-training. We use the ‘Winter21 variant of ImageNet-21K-P’ (refered to as ImageNet-21K) dataset ridnik2021imagenet 111downloaded from: https://www.image-net.org/download-images.php for large-scale pre-training, which contains 11060223 training images and 522500 testing images from 10450 classes. Only the train set was used. We pre-train iLLaMA-B/L on ImageNet-21K for 90 epochs using a constant soft mask schedule, with cutoff epochs and learning rate warmup epochs set to 30 and 5, respectively. Detailed recipes can be found in Appendix B.2.
ImageNet-1K fine-tuning. We fine-tune the iLLaMA-B model, trained on ImageNet-1K, at a resolution of 384$\times$384. Similarly, the iLLaMA-B/L model, trained on ImageNet-21K, was fine-tuned at resolutions of 224$\times$224 and 384$\times$384, respectively. All fine-tuning was conducted for 30 epochs using the AdamW optimizer. We follow DeiT touvron2021training for interpolating positional embeddings to allow our iLLaMA to handle inputs at a higher resolution. Other settings are detailed in Appendix B.3.
@@ 4.3 Model Calibration and Shape-Texture Bias
Model calibration. Model calibration represents the relationship between a model’s precision and confidence across samples of varying difficulty, i.e., poor-calibrated models tend to produce overly confident yet incorrect predictions, whereas well-calibrated models demonstrate a strong correlation between confidence and accuracy guo2017calibration . Calibration is commonly measured using the Expected Calibration Error (ECE), where a lower ECE is favorable. ECE results for different models on ImageNet-1K are presented in Table 8. The calibration of iLLaMA is lower than that of DeiT3, suggesting that the output confidence of iLLaMA is more reliable. We also plot the reliability diagrams vishniakov2023convnet to intuitively compare the calibration of different models, detailed in Appendix F.
Shape-texture bias. Shape-texture bias measures the extent to which the model relies on the shape or texture of the image when performing recognition geirhos2018imagenet . We generally prefer models to mimic human eye behavior, relying more on shape rather than texture tuli2021convolutional ; geirhos2020shortcut . We calculate the shape ratio for all models on cue-conflict images and report the results in Table 8, following vishniakov2023convnet . Our iLLaMA shows the largest shape ratio of $41.45\%$ among the three compared baselines, suggesting the potential of the LLM architecture for vision. Detailed results can be found in Appendix G.
@@ 4.4 Compatibility with Quantization
Since a practical goal for neural networks is deployment on low-bit hardware chips, we further examine iLLaMA’s compatibility with quantization. We basically follow Q-ViT li2022q to apply quantization-aware training (QAT) to iLLaMA, with weights and activations of all blocks’ FFN and causal self-attention layers to 8 bits. Quantization recipes and results are shown in Appendix B.4 and Table 8, respectively. Different sizes of low-bit iLLaMA maintain accuracy well, and 8-bit iLLaMA-T is even compete favorably with the full-precision DeiT-T touvron2021training ($72.4\%$ v.s. $72.2\%$).
@@ 4.5 Transferability on Downstream Tasks
CIFAR transfer learning. We fine-tune ViT-T and iLLaMA-T on the CIFAR datasets krizhevsky2009learning , including an ablation of the soft mask on iLLaMA. Detailed recipes are provided in Appendix B.5. iLLaMA’s performance on CIFAR datasets is essentially on par with ViT, assuring that iLLaMA can be confidently applied in the transfer learning field as a practical alternative to ViT. Additionally, soft mask is helpful in the relatively complicated CIFAR100, demonstrating its generalizability.
@@ 5 Conclusions
In the general trend of using encoder to build vision architectures, the suitability for decoder-only architecture still exists a research gap. We systematically studies whether decoder, an architecture that has shown amazing potential in LLMs, can also take root in learning visual representation through visual supervised training. The key component – causal self-attention we used – is not novel and is inherited from existing LLM architectures, but we propose pivotal techniques, i.e., PS [cls] and soft mask, to effectively adapt them to visual tasks. We hope that this work will inspire more exploration towards a generic pure decoder architecture that fully unifies vision and text.
Appendix
In this Appendix, we detail our network configuration (§A), outline experimental recipe to replicate our study (§B), share insights from the iLLaMA exploration (§C), , and analyze both the causal self-attention (§D) and soft mask techniques (§E). Additionally, we explore iLLaMA’s calibration property (§F), shape-texture bias (§G), loss landscape (§H), its limitations (§I), and its potential societal impacts (§J).
@@ Appendix A Network Configuration
Appendix B Experimental Settings
B.1 Training Recipe in Section 3
@@
B.2 ImageNet (Pre-)training Recipe
@@
For the iLLaMA-T/S/B models, we train directly on ImageNet-1K and discover that models of different sizes are suited to different soft mask settings. For instance, the soft mask schedules are set to constant/linear/linear, respectively, with cutoff epochs designated as 50/50/25.
We pre-trained the iLLaMA-B/L models on ImageNet-21K for 90 epochs, adhering to the practices in liu2022convnet . We set the cutoff epochs to 30, indicating that the iLLaMA models’ self-attention fully transitions to causal self-attention after 30 epochs.
@@
B.3 ImageNet Fine-tuning Recipe
@@
Additionally, we standardized the cutoff epoch at 0 for our ImageNet-1K fine-tuning experiments, ensuring the application of a causal mask in self-attention to align with the LLaMA architecture. We also opted not to use learning rate warmup.
@@
B.4 Quantization-aware Training Recipe
@@
B.5 CIFAR Transfer Learning Recipe
@@
Appendix C Designing iLLaMA: detailed results
@@
Appendix D Rank Analysis of causal Self-attention
@@
Appendix E Analysis for Soft Mask Method
@@
Appendix F Model Calibration
@@
Appendix G Shape-Texture Bias
@@
Appendix H Loss Landscape
@@ Appendix I Limitations
Appendix J Societal Impact

### v4

@@ 3.7 Analysis of causal Self-attention
Attention map rank. We examine the representation learning power of causal attention through a spectrum analysis. Following wang2020linformer ; shu2021adder , we perform singular value decomposition on the attention maps of the pre-trained ViT-T/16 and iLLaMA-T/16 models. Next, we sort the singular values and plot a curve illustrating the relationship between the cumulative normalized singular values and matrix indices. The results are conducted using $30$ images randomly selected from the ImageNet-1K validation set. As shown in Figure 4, the curve of ViT exhibits concave function characteristics, while the curve of iLLaMA is close to a linear function, indicating a more uniform distribution of singular values in iLLaMA’s attention map. Approximating the matrix rank by the index at which the cumulative normalized singular value reaches 0.8, we observe that the index value of iLLaMA is about 48 higher than that of ViT ($\sim$129-th v.s. $\sim$81-th). Under such premise, compared to ViT, the attention map of iLLaMA can be approximated with a certain error by a higher-rank matrix. Accordingly, the rank of the attention map may affect the expressive capabilities of the learned representations dong2021attention , suggesting that the causal self-attention in iLLaMA has the potential to learn complex visual representations, as empirically demonstrated in Section 4.2. Detailed results for different layers are provided in Figure 11 of Appendix E.
@@ 4.1 Ablation Study
Influence of soft mask for training and testing loss. Deep neural networks often face underfitting, marked by difficulty in continuously reducing training loss and resulting in poor test accuracy liu2023dropout . We compare the training and testing losses of the iLLaMA-T/16 and iLLaMA-B/16 models with and without the use of the soft mask strategy. As shown in Table 4, soft mask can reduce training loss in both regimes, mitigating potential underfitting issue and reducing testing loss.
@@ 4.2 Comparison with Recent Architectures on ImageNet-1K Classification
We conducted experiments on the ImageNet-1K deng2009imagenet benchmark with different model sizes (i.e., iLLaMA-T/S/B/L). Detailed architecture configurations are shown in Table 11 of Appendix A. Our ImageNet-1K/21K (pre-)training and ImageNet-1K fine-tuning recipes are shown in Table 13 and Table 14 of Appendix C. We also study the use of LLaMA2-7B pre-trained weights for iLLaMA initialization, and the paradigm and results are detailed in Figure 9 and Table 18 in Appendix I.
ImageNet-1K training. We train iLLaMA-T/S/B on ImageNet-1K for 300 epochs with AdamW optimizer loshchilov2017decoupled and a batch size of 4096. The ImageNet-1K trained iLLaMA-T/B models are, in fact, the outcome of the explorations completed in Section 3.6. For the settings of soft mask schedule, cutoff epochs, and learning rate warmup epochs, we tune slightly for the iLLaMA-S model.
ImageNet-21K pre-training. We use the ‘Winter21 variant of ImageNet-21K-P’ (refered to as ImageNet-21K) dataset ridnik2021imagenet 111downloaded from: https://www.image-net.org/download-images.php for the large-scale pre-training of our iLLaMA, which contains 11,060,223 training images and 522,500 testing images from 10,450 classes. Only the training set was used. We pre-train iLLaMA-B/L on ImageNet-21K for 90 epochs using a constant soft mask schedule, with cutoff epochs and learning rate warmup epochs set to 30 and 5, respectively.
ImageNet-1K fine-tuning. For iLLaMA-B model trained on ImageNet-1K, we fine-tune at a resolution of 384$\times$384. Similarly, for the iLLaMA-B/L model trained on ImageNet-21K, we fine-tune at resolutions of 224$\times$224 and 384$\times$384, respectively. All fine-tuning was conducted for 30 epochs using the AdamW optimizer. We follow DeiT touvron2021training for interpolating positional embeddings to allow our iLLaMA to handle inputs at a higher resolution.
@@ 4.3 Model Calibration and Shape-Texture Bias
Model calibration. Model calibration represents the relationship between a model’s precision and confidence across samples of varying difficulty, i.e., poor-calibrated models tend to produce overly confident yet incorrect predictions, whereas well-calibrated models demonstrate a strong correlation between confidence and accuracy guo2017calibration . Calibration is commonly measured using the Expected Calibration Error (ECE), where a lower ECE is favorable. ECE results for different models on ImageNet-1K are presented in Table 8. The calibration of iLLaMA is lower than that of DeiT3, suggesting a more reliable output confidence. We also plot the reliability diagrams vishniakov2023convnet to intuitively compare the calibration of different models, as shown in Figure 7 of Appendix G.
Shape-texture bias. Shape-texture bias measures the extent to which the model relies on the shape or texture of the image when performing recognition geirhos2018imagenet . We generally prefer models to mimic human eye behavior, relying more on shape rather than texture tuli2021convolutional ; geirhos2020shortcut . We calculate the shape ratio for all models on cue-conflict images and report the results in Table 8, following vishniakov2023convnet . Our iLLaMA shows the largest shape ratio of $41.45\%$ among the three compared baselines, suggesting the potential of the LLM architecture for vision. Detailed results are provided in Figure 8 of Appendix H.
@@ 4.4 Compatibility with Quantization
Since a practical goal for neural networks is deployment on low-bit hardware chips, we further examine iLLaMA’s compatibility with quantization. We basically follow Q-ViT li2022q to apply quantization-aware training (QAT) to iLLaMA, with weights and activations of all blocks’ FFN and causal self-attention layers to 8 bits. Quantization recipes and results are shown in Table 15 of Appendix C.4 and Table 8. Different sizes of low-bit iLLaMA maintain accuracy well, and 8-bit iLLaMA-T is even compete favorably with the full-precision DeiT-T touvron2021training ($72.4\%$ v.s. $72.2\%$).
@@ 4.5 Transferability on Downstream Tasks
CIFAR transfer learning. We fine-tune ViT-T and iLLaMA-T on the CIFAR datasets krizhevsky2009learning , including an ablation of the soft mask on iLLaMA. Detailed recipes are shown in Table 16 of Appendix C.5. iLLaMA’s performance on CIFAR datasets is essentially on par with ViT, assuring that iLLaMA can be confidently applied in the transfer learning field as a practical alternative to ViT. Additionally, soft mask is helpful in the relatively complicated CIFAR100, demonstrating its generalizability.
@@ 5 Conclusions
In this paper, we systematically studies whether Transformer decoder, an architecture that has shown amazing potential in LLMs, can also take root in learning visual representation through straightforward supervised training. The key component – causal self-attention we used – is not novel and is inherited from existing LLM architectures, but we propose pivotal techniques, i.e., PS [cls] and soft mask strategies, to effectively adapt them to visual tasks. The proposed iLLaMA outperforms many ConvNets, ViTs, and MLPs on imagenet, and demonstrates robust quantization compatibility, calibration, and shape-texture bias, thereby showing its practicality. We hope that this work will inspire a rethinking of generic yet practical architecture that can fully unify both vision and text.
@@ Appendix A Network Configuration
Appendix B PyTorch-like Code of iLLaMA Causal Self-attention
The PyTorch-like implementation of our iLLaMA causal self-attention is shown as Algorithm 1. The iLLaMA code exhibits a high degree of similarity in structure and composition to the official LLaMA code  [^2] https://github.com/meta-llama/llama released by Meta, potentially offering considerable coding cost savings in developing a unified vision and language network with such architecture.
Algorithm 1 PyTorch code of iLLaMA causal self-attention
import torch
import torch.nn as nn
def reshape_for_broadcast(freqs_cis: torch.Tensor, x: torch.Tensor):
ndim = x.ndim
assert 0 <= 1 < ndim
assert freqs_cis.shape == (x.shape[1], x.shape[-1])
shape = [d if i == 1 or i == ndim - 1 else 1 for i, d in enumerate(x.shape)]
return freqs_cis.view(*shape)
def apply_rotary_emb(
xq: torch.Tensor,
xk: torch.Tensor,
freqs_cis: torch.Tensor,
) -> Tuple[torch.Tensor, torch.Tensor]:
xq_ = torch.view_as_complex(xq.float().reshape(*xq.shape[:-1], -1, 2))
xk_ = torch.view_as_complex(xk.float().reshape(*xk.shape[:-1], -1, 2))
freqs_cis = reshape_for_broadcast(freqs_cis, xq_)
xq_out = torch.view_as_real(xq_ * freqs_cis).flatten(3)
xk_out = torch.view_as_real(xk_ * freqs_cis).flatten(3)
return xq_out.type_as(xq), xk_out.type_as(xk)
class Attention(nn.Module):
def __init__(self, dim, num_heads=8, qkv_bias=False, qk_scale=None, attn_drop=0., proj_drop=0.):
super().__init__()
self.num_heads = num_heads
head_dim = dim // num_heads
# NOTE scale factor was wrong in my original version, can set manually to be compat with prev weights
self.scale = qk_scale or head_dim ** -0.5
self.qkv = nn.Linear(dim, dim * 3, bias=qkv_bias)
self.proj = nn.Linear(dim, dim)
def forward(self, x: torch.Tensor, freqs_cis: torch.Tensor, mask: Optional[torch.Tensor]):
B, N, C = x.shape
qkv = self.qkv(x).reshape(B, N, 3, self.num_heads, C // self.num_heads).permute(2, 0, 1, 3, 4) # [3, B, N, self.num_heads, C // self.num_heads]
q, k, v = qkv[0], qkv[1], qkv[2] # make torchscript happy (cannot use tensor as tuple) # [B, N, self.num_heads, C // self.num_heads]
q, k = apply_rotary_emb(q, k, freqs_cis=freqs_cis)
q = q.transpose(1, 2) # [B, self.num_heads, N, C // self.num_heads]
k = k.transpose(1, 2) # [B, self.num_heads, N, C // self.num_heads]
v = v.transpose(1, 2) # [B, self.num_heads, N, C // self.num_heads]
attn = (q @ k.transpose(-2, -1)) * self.scale # [B, self.num_heads, N, N]
attn = attn.softmax(dim=-1)
if mask is not None:
attn = attn * mask # (B, H, N, N)
x = (attn @ v).transpose(1, 2).reshape(B, N, C)
x = self.proj(x)
return x
Appendix C Experimental Settings
C.1 Training Recipe in Section 3
@@
C.2 ImageNet (Pre-)training Recipe
@@
For the iLLaMA-T/S/B models, we train directly on ImageNet-1K and discover that models of different sizes are suited to different soft mask settings. For instance, the soft mask schedules are set to constant/linear/linear, respectively, with cutoff epochs designated as 50/50/25. We train the iLLaMA-T/S/B models using 8 A100 GPUs.
We pre-trained the iLLaMA-B/L models on ImageNet-21K for 90 epochs, adhering to the practices in liu2022convnet . We set the cutoff epochs to 30, indicating that the iLLaMA models’ self-attention fully transitions to causal self-attention after 30 epochs. We pre-train the iLLaMA-B/L models using 8 A100 GPUs.
@@
C.3 ImageNet Fine-tuning Recipe
@@
Additionally, we standardized the cutoff epoch at 0 for our ImageNet-1K fine-tuning experiments, ensuring the application of a causal mask in self-attention to align with the LLaMA architecture. We also opted not to use learning rate warmup. We fine-tune the models using 8 A100 GPUs.
@@
C.4 Quantization-aware Training Recipe
@@
C.5 CIFAR Transfer Learning Recipe
@@
Appendix D Designing iLLaMA: detailed results
@@
Appendix E Rank Analysis of causal Self-attention
@@
Appendix F Analysis for Soft Mask Method
@@
Appendix G Model Calibration
@@
Appendix H Shape-Texture Bias
@@
Appendix I Initializating iLLaMA Using Pre-trained LLaMA
Previous studies zhang2024multimodal have demonstrated that data unrelated to the image modality can be used to improve the performance of visual models. In fact, the pre-training dataset of LLaMA, which is entirely text, is irrelevant to the visual tasks that iLLaMA addresses. More important, the architectural components of iLLaMA are aligned with those of LLaMA. This alignment facilitates our exploration of using LLaMA’s parameters to initialize iLLaMA, allowing us to fully exploit the potential of the weights of pre-trained LLMs.
Table 18: Ablation results of weight selection of iLLaMA using LLaMA2-7B pre-trained weights.
Model Initialization Tiny Small Base
iLLaMA w/ weight selection 74.5 79.9 81.4
iLLaMA w/o weight selection 75.0 79.9 81.6
We use the pre-trained LLaMA2-7B touvron2023llama2 to initialize our iLLaMA, instead of training from scratch. To match the dimensions of the weights, we employ the weight selection xu2023initializing method to initialize iLLaMA-T/S/B using a subset of the LLaMA2-7B pre-trained weights. Next, we train and evaluate the iLLaMA models, which are initialized using LLaMA2-7B, on the ImageNet-1K dataset. Other hyperparameter settings are consistent with Section 4.2. The results are shown in Table 18. Using LLaMA2 to initialize iLLaMA does not yield significant performance improvements. We attribute this to two main reasons: 1) The size difference between the two models is considerably large (LLaMA-2-7B’s 7B parameters vs. iLLaMA-T’s 5.7M parameters), resulting in a limited proportion of selected weights compared to meaningful pre-trained weights. 2) The training strategy was not adequately optimized. We believe that when using parameter inheritance, the corresponding training strategy should also be adjusted. However, we continued to use the training recipe designed for training from scratch.
Figure 9: iLLaMA initialization by pre-trained LLaMA2-7B touvron2023llama2 using weight selection xu2023initializing .
Appendix J Loss Landscape
@@ Appendix I Initializating iLLaMA Using Pre-trained LLaMA
Appendix L Societal Impact